OpenStudy (anonymous):

5/22 in simpiliest form

OpenStudy (paki):

5/22 = 0.2272

OpenStudy (anonymous):

5/22 is in its simplest form, unless they want a decimal which would be what Paki replied.

OpenStudy (anonymous):

1/2 x 5/11 =5/22 now need the simpiliest form

OpenStudy (anonymous):

Paki gave you the answer.